# On the new and accurate (Goudsmit-Saunderson) model for describing e-/e+ multiple Coulomb scattering

### v1

@@ A.1 Goudsmit-Saunderson series
$\int_{-1}^{1}P_{\ell}(x)P_{k}(x)\mathrm{d}x=\left\{\begin{array}[]{l l}C_{\ell}=\frac{2}{2\ell+1}&\text{if $\ell=k$ }\\ 0&\text{if $\ell\neq k$ }\\ \end{array}\right.$ (A.1)
@@
$\begin{split}f_{n=2}(\cos(\theta))=&\int_{-1}^{+1}\int_{0}^{2\pi}f_{1}(\cos(\theta_{1}))f_{1}(\cos(\theta_{2}))\mathrm{d}\phi_{1}\mathrm{d}(\cos(\theta_{1}))\\ &=\int_{-1}^{+1}\int_{0}^{2\pi}\left[\sum_{\ell=0}^{\infty}\frac{2\ell+1}{4\pi}F_{\ell}P_{\ell}(\cos(\theta_{1}))\right]\left[\sum_{\ell^{\prime}=0}^{\infty}\frac{2\ell^{\prime}+1}{4\pi}F_{\ell^{\prime}}P_{\ell^{\prime}}(\cos(\theta_{2}))\right]\mathrm{d}\phi_{1}\mathrm{d}(\cos(\theta_{1}))\end{split}$ (A.10)
@@ A.4 Optimal parameter of the variable transformation
$\int_{-1}^{+1}P_{\ell}(x)P_{k}(x)x^{2}\mathrm{d}x=\left\{\begin{array}[]{ll}\frac{2(\ell+1)(\ell+2)}{(2\ell+1)(2\ell+3)(2\ell+5)}&\quad\mathrm{if}\;k=\ell+2\\ &\\ \frac{2(2\ell^{2}+2\ell-1)}{(2\ell-1)(2\ell+1)(2\ell+3)}&\quad\mathrm{if}\;k=\ell\\ &\\ \frac{2\ell(\ell-1)}{(2\ell-3)(2\ell-1)(2\ell+1)}&\quad\mathrm{if}\;k=\ell-2\\ \end{array}\right.$ (A.49)
@@
$G_{1}s/\lambda$ $A$ $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@
$G_{1}s/\lambda$ $A$ $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@
$G_{1}s/\lambda$ $A$ $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@
$G_{1}s/\lambda$ $A$ $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@ A.5 On the run time interpolation of the pre-computed PDF-s
Let $x\in\mathbb{R}\;|\;x\in[x_{\text{min}},x_{\text{max}}],x_{\text{max}}>x_{\text{min}}$ a continus variable. Dividing the $[x_{\text{min}},x_{\text{max}}]$ interval into $N$ equal sub-intervals (bins) corresponds to introduce the $\{x_{0},x_{1},...,x_{N}\}\;N+1$ discrete values such that
@@
Then the index $k$, such that $x_{k}\leq x<x_{k+1}$ for any $x\in\mathbb{R}\;|\;x\in[x_{\text{min}},x_{\text{max}})$ continuous value, is the $k=[(x-x_{0})/\Delta]$ where []delimited-[][$\cdot$][ ⋅ ] is used for the integer part (or floor) of the expression. Furthermore, let $y(x)$ a single value continuous function of $x$, i.e. $y:[x_{\text{min}},x_{\text{max}}]\to\mathbb{R}$. Linear interpolation of $y(x)$ at any $x\in[x_{\text{min}},x_{\text{max}})$ such that $x_{k}\leq x<x_{k+1}$ corresponds to
@@
that shows more explicitly that the result of the linear interpolation $\tilde{y}(x)$ is nothing but the linear combination of the two $y_{k},\;y_{k+1}$ exact function values taken at $x_{k}$ and $x_{k+1}$ with the $1-w$ and $w\equiv(x-x_{k})/(x_{k+1}-x_{k})$ weights respectively.
@@
This shows clearly that the result of the linear interpolation of the pdf $\tilde{p}(u;x)$ is nothing but a mixture distribution of the exact $p(u;x_{k})$ and $p(u;x_{k+1})$ pdf-s with the $1-w$ and $w$ weights respectively. This is because the weights $1-w$ and $w$ are both $\geq 0$ while sum up to unity. Therefore, providing $u$ samples distributed according to the $\tilde{p}(u;x)$ interpolated pdf can be done by generating samples form $p(u;x_{k})$ with a probability of $1-w$ and from $p(u;x_{k+1})$ with a probability of $w\equiv(x-x_{k})/(x_{k+1}-x_{k})$.
@@
    generate a random $\begin{array}[]{ll}\text{generate a random number:}&\eta\in\mathcal{U}(0,1)\\ \text{sample }u\text{ from }p(u;x_{k+1})&\text{if }\eta<(x-x_{k})/(x_{k+1}-x_{k}):\\ \text{sample }u\text{ from }p(u;x_{k})&\text{otherwise}\\ \end{array}$ (A.60)

### v2

@@ 4 Electron-step algorithm
The algorithm implemented for calculating the post-step point in a condensed history (CH) simulation step is described in this section. The implemented algorithm is the one developed by Kawrakow in [6] including the energy loss correction from [2]. Only the essentials required for the calculations are summarised in this section below while all details can be found in [6, 2].
[…]
In order to reduce the approximation error, the CH step is divided to two equal sub-steps in the algorithm proposed in [6]. Polar and azimuthal scattering angles are sampled for the first $\theta_{1},\phi_{1}$ and second $\theta_{2},\phi_{2}$ sub-steps using the pre-computed GS angular distributions for $\theta_{1}$ and $\theta_{2}$ as described in section 2.4.3. Having a coordinate system with $z$-axis pointing to the particle momentum direction, i.e. an initial momentum direction $\mathbf{\hat{d}_{0}}=(0,0,1)$ in this system, the direction after the first scattering
$\mathbf{\hat{d}_{1}}=R_{z}(\phi_{1})\,R_{y}(\theta_{1})\,\mathbf{\hat{d}_{0}}=\begin{pmatrix}\sin\theta_{1}\cos\phi_{1}\\[2.0pt] \sin\theta_{1}\sin\phi_{1}\\[2.0pt] \cos\theta_{1}\end{pmatrix}\equiv\begin{pmatrix}u_{1}\\[2.0pt] v_{1}\\[2.0pt] w_{1}\end{pmatrix}$. (82)
where $R_{z},R_{y}$ are the matrices for rotation around the $z$ and $y$ axes. The final direction, i.e. after the second scattering, then
$\mathbf{\hat{d}_{f}}$ $=\begin{pmatrix}\sin\theta_{1}\cos\phi_{1}\,\cos\theta_{2}\cos\phi_{2}-\sin\theta_{1}\sin\phi_{1}\,\sin\phi_{2}+\cos\theta_{1}\,\sin\theta_{2}\cos\phi_{2}\\[6.0pt] \sin\theta_{1}\cos\phi_{1}\,\cos\theta_{2}\sin\phi_{2}+\sin\theta_{1}\sin\phi_{1}\,\cos\phi_{2}+\cos\theta_{1}\,\sin\theta_{2}\sin\phi_{2}\\[6.0pt] -\sin\theta_{1}\cos\phi_{1}\,\sin\theta_{2}+\cos\theta_{1}\,\cos\theta_{2}\end{pmatrix}$ (83)
$=\begin{pmatrix}u_{1}\,\cos\theta_{2}\cos\phi_{2}-v_{1}\,\sin\phi_{2}+\cos\theta_{1}\,\sin\theta_{2}\cos\phi_{2}\\[6.0pt] u_{1}\,\cos\theta_{2}\sin\phi_{2}+v_{1}\,\cos\phi_{2}+\cos\theta_{1}\,\sin\theta_{2}\sin\phi_{2}\\[6.0pt] -u_{1}\,\sin\theta_{2}+\cos\theta_{1}\,\cos\theta_{2}\end{pmatrix}$ (84)
$=\begin{pmatrix}\cos\phi_{2}\,t_{m}-v_{1}\,\sin\phi_{2}\\[6.0pt] \sin\phi_{2}\,t_{m}+v_{1}\,\cos\phi_{2}\\[6.0pt] -u_{1}\,\sin\theta_{2}+\cos\theta_{1}\,\cos\theta_{2}\end{pmatrix}\;\equiv\;\begin{pmatrix}\sin\theta\,\cos\phi\\[6.0pt] \sin\theta\,\sin\phi\\[6.0pt] \cos\theta\end{pmatrix}\equiv\begin{pmatrix}u_{f}\\[2.0pt] v_{f}\\[2.0pt] w_{f}\end{pmatrix}$. (85)
where $t_{m}\;=\;\sin\theta_{1}\cos\phi_{1}\,\cos\theta_{2}+\cos\theta_{1}\,\sin\theta_{2}$ while $\theta,\phi$ are the combined polar and azimuthal scattering angles. Note that $\theta_{i}$ and $\phi_{i}$ are samples from the same distributions therefore the indices $1\to 2$ are interchangeable. Having the pre-step point in the origin of the coordinate system, the post-step point coordinates of the CH step might be given [6] as
$x_{f}$ $=s\Big[\,\eta_{1}\sin\theta_{1}\cos\phi_{1}$ $+$ $\eta_{2}\sin\theta_{2}\big(\cos\phi_{1}\cos\phi_{2}-\cos\theta_{1}\sin\phi_{1}\sin\phi_{2}\big)$ $+$ $(1-\eta_{0}-\eta_{1}-\eta_{2})\sin\theta\cos\phi\,\Big],$ (86)
$y_{f}$ $=s\Big[\,\eta_{1}\sin\theta_{1}\sin\phi_{1}$ $+$ $\eta_{2}\sin\theta_{2}\big(\sin\phi_{1}\cos\phi_{2}+\cos\theta_{1}\cos\phi_{1}\sin\phi_{2}\big)$ $+$ $(1-\eta_{0}-\eta_{1}-\eta_{2})\sin\theta\sin\phi\,\Big],$
$z_{f}$ $=s\Big[\,\eta_{0}+\eta_{1}\cos\theta_{1}$ $+$ $\eta_{2}\cos\theta_{2}$ $+$ $(1-\eta_{0}-\eta_{1}-\eta_{2})\cos\theta\,\Big].$
The coefficients $\eta_{0},\eta_{1},\eta_{2}$ were then determined such that the moments considered are well reproduced leading to the following equations
$\eta_{0}$ $=(1-\eta)/2,$ (87)
$\eta_{1}$ $=\eta\delta,$
$\eta_{2}$ $=\eta(1-\delta)$
with $\eta\in[0,1]$ sampled from the $2\eta\mathrm{d}\eta$ distribution and
$\delta=\frac{1}{2}+\frac{\sqrt{6}}{6}-\left(\frac{1}{4\sqrt{6}}-\gamma\frac{4-\sqrt{6}}{24\sqrt{6}}\right)Q_{1}$ (88)
with the
$Q_{1}$ $\equiv sG_{\ell=1}/\lambda$ with $G_{\ell=1}^{SR}=2A\left[\ln\left(1+\frac{1}{A}\right)(A+1)-1\right],$ (89)
$\gamma$ $\equiv G_{\ell=2}/G_{\ell=1}$ with $G_{\ell=2}^{SR}=6A(1+A)\left[(1+2A)\ln\left(1+\frac{1}{A}\right)-2\right]$
notation introduced in section 3.3 for these quantities that determines the first and second moments of the GS angular distribution along a CH step with a path-length of $s$. The expression for $G_{\ell=1}$ and $G_{\ell=2}$ are given by Eq.(A.40) and Eq.(A.42) as derived in APPENDIX A.3 under the Screened Rutherford DCS approximation with a screening parameter of $A$. Introducing $a_{1}=(1-\eta_{0}-\eta_{1}-\eta_{2})$ and $u_{2}=\sin\theta_{2}\cos\phi_{2},\;v_{2}=\sin\theta_{2}\sin\phi_{2}$ the post-step coordinates can be written in the compact form
$x_{f}$ $=s\Big[\,\eta_{1}u_{1}$ $+$ $\eta_{2}\big(\cos\phi_{1}u_{2}-\cos\theta_{1}\sin\phi_{1}v_{2}\big)$ $+$ $a_{1}u_{f}\,\Big],$ (90)
$y_{f}$ $=s\Big[\,\eta_{1}v_{1}$ $+$ $\eta_{2}\big(\sin\phi_{1}u_{2}+\cos\theta_{1}\cos\phi_{1}v_{2}\big)$ $+$ $a_{1}v_{f}\,\Big],$
$z_{f}$ $=s\Big[\,\eta_{0}+\eta_{1}\cos\theta_{1}$ $+$ $\eta_{2}\cos\theta_{2}$ $+$ $a_{1}w_{f}\,\Big].$
With the above expressions for $\eta_{0},\eta_{1},\eta_{2}$, now $a_{1}=(1-\eta_{0}-\eta_{1}-\eta_{2})=\eta_{0}$.
The above algorithm for computing the CH post-step point position was obtained to reproduce the moments of the spatial distributions and spatial-angular correlations of the Lewis’ theory when neglecting the energy loss along the step. Corrections for accounting energy loss were derived in [2] (see Appendix B there) by comparing the corresponding first and second spatial moments, i.e. the mean longitudinal <z><z> and squared lateral $r^{2}>=<x^{2}+y^{2}>$ displacements, to the results of Lewis’ theory including the energy loss. The first order correction (in $\Delta E$) corresponds to changing the parameters $\eta_{0}\to eta_{0}(1+\alpha)$ and $\delta\to\delta+\alpha_{2}$ with the following expressions for $\alpha_{1},\alpha_{2}$ resulting from the above-mentioned comparison [2]
$\alpha_{1}$ $=-\frac{\kappa_{1}^{\prime}(\tilde{E})}{\kappa_{1}(\tilde{E})}\frac{\Delta E}{2}+\mathcal{O}(\Delta E^{2}),$ (91)
$\alpha_{2}$ $=\left[\frac{\kappa_{2}^{\prime}(\tilde{E})}{\kappa_{2}(\tilde{E})}-\frac{\kappa_{1}^{\prime}(\tilde{E})}{\kappa_{1}(\tilde{E})}\right]\frac{\Delta E}{2\sqrt{6}}+\mathcal{O}(\Delta E^{2})$
with $\kappa_{\ell}\equiv G_{\ell}/\lambda=\lambda_{\ell}^{-1}=\Sigma_{\ell}$ being the $\ell$-th macroscopic transport cross section (or inverse transport mean free path) as given by the definition Eq.(7), $\tilde{E}=E_{0}-\Delta E/2$ is the mid-point energy along the step as used in section 3.1 while $(\cdot)^{\prime}$ denotes their derivatives with respect to $E$. The expressions for $\Sigma_{\ell=1}$ and $\Sigma_{\ell=2}$ as given in [2] Eqs.(A5) (denoted by $\kappa_{\ell=1},\kappa_{\ell=2}$ there) can be obtained by starting the general definition $\Sigma_{\ell}\equiv\lambda_{\ell}^{-1}=G_{\ell}/\lambda$ as given by Eq.(7) then substituting the SR DCS forms of $G_{\ell=1}$ and $G_{ell=2}$ as given by Eqs.(A.40,A.42) with the expression for $\lambda$ as given by Eq.(66) (with the usual approximation of dropping the $1/(1+A)$ factor from this latter)
$\Sigma_{\ell=1}^{(SR)}$ $=2A\frac{b_{c}}{\beta^{2}}\left[\ln\left(1+\frac{1}{A}\right)(1+A)-1\right],$ (92)
$\Sigma_{\ell=2}^{(SR)}$ $=6A\frac{b_{c}}{\beta^{2}}(1+A)\left[(1+2A)\ln\left(1+\frac{1}{A}\right)-2\right]$
which are the same as Eqs.(A5) in [2] when substituting Eq.(68) for $A$. Substituting these into the above expressions for $\alpha_{1}$ and $\alpha_{2}$ yields [2]999the $\alpha_{1}^{2}/(4\sqrt{6})$ leading 2nd order contribution is also added to $\alpha_{2}$ in the EGSnrc implementation
$\alpha_{1}$ $\approx\left[\frac{2+2\tilde{\tau}+\tilde{\tau}^{2}}{1+\tilde{\tau}}-\frac{1+\tilde{\tau}}{\ln\!\big(1+1/\tilde{A}\big)\,(1+\tilde{A})-1}\right]\,\frac{\tilde{\epsilon}}{2+\tilde{\tau}}\,$, (93)
$\alpha_{2}$ $\approx\frac{\tilde{\epsilon}}{\sqrt{6}\,(1+\tilde{\tau})(2+\tilde{\tau})\,\big[\ln\!\big(1+1/\tilde{A}\big)\,(1+\tilde{A})-1\big]\,\big[\ln\!\big(1+1/\tilde{A}\big)\,(1+2\tilde{A})-2\big]}\,.$
where $\tilde{\epsilon}=\Delta E/\tilde{E}$, $\tilde{\tau}=\tilde{E}/m_{e}$ is the mid-point energy in electron rest-mass $m_{e}$ units just as in Section 3.1 and $\tilde{A}$ is the screening parameter evaluated at the mid-point energy of the step (i.e. taking $(pc)^{2}=\tilde{E}(\tilde{E}+2m_{e})$ in Eq.(68)).
When accounting energy loss along the CH step, $\eta$ is sampled from the $2\eta\mathrm{d}\eta$ distribution, $\gamma$ and $\delta$ are computed by Eqs.(89,88), $\eta_{0},\eta_{1},\eta_{2}$ can be calculated by Eq.(87) then $\eta_{0}=\eta_{0}(1+\alpha_{1})$ and $\delta=\delta+\alpha_{2}$ can be obtained after evaluating the $\alpha_{1}$ and $\alpha_{2}$ energy loss correction parameters given by Eqs.(93). With these forms of $\eta_{0}$ and $\delta$, the parameter $a_{1}$, introduced at Eq.(90), have the $a_{1}\equiv 1-\eta_{0}-\eta_{1}-\eta_{2}=\eta_{0}(1-\alpha_{1})$ form now. The post-step position of the CH step can be calculated then by using these parameters in Eq.(90) after dividing the step into two equal sub-steps, generating $\theta_{1},\phi_{1}$ and $\theta_{2},\phi_{2}$ for the sub-steps and calculating the corresponding final direction components as given by Eq.(86).
[…]
5 Summary and final notes
@@ A.1 Goudsmit-Saunderson series
[…]
@@
$\begin{split}f_{n=2}(\cos(\theta))=&\int_{-1}^{+1}\int_{0}^{2\pi}f_{1}(\cos(\theta_{1}))f_{1}(\cos(\theta_{2}))\mathrm{d}\phi_{1}\mathrm{d}(\cos(\theta_{1}))\\ &\hskip-56.9055pt=\int_{-1}^{+1}\int_{0}^{2\pi}\left[\sum_{\ell=0}^{\infty}\frac{2\ell+1}{4\pi}F_{\ell}P_{\ell}(\cos(\theta_{1}))\right]\left[\sum_{\ell^{\prime}=0}^{\infty}\frac{2\ell^{\prime}+1}{4\pi}F_{\ell^{\prime}}P_{\ell^{\prime}}(\cos(\theta_{2}))\right]\mathrm{d}\phi_{1}\mathrm{d}(\cos(\theta_{1}))\end{split}$ (A.10)
@@ A.4 Optimal parameter of the variable transformation
[…]
@@
$G_{1}s/\lambda$ AA $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@
$G_{1}s/\lambda$ AA $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@
$G_{1}s/\lambda$ AA $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@
$G_{1}s/\lambda$ AA $a$ $w^{2}=a/A$ $\tilde{a}=\tilde{w}^{2}A$
@@ A.5 On the run time interpolation of the pre-computed PDF-s
Let $x\in\mathbb{R}\;|\;x\in[x_{\text{min}},x_{\text{max}}],x_{\text{max}}>x_{\text{min}}$ a continus variable. Dividing the $x_{\text{min}},x_{\text{max}}]$ interval into $N$ equal sub-intervals (bins) corresponds to introduce the $\{x_{0},x_{1},...,x_{N}\}\;N+1$ discrete values such that
@@
Then the index $k$, such that $x_{k}\leq x<x_{k+1}$ for any $x\in\mathbb{R}\;|\;x\in[x_{\text{min}},x_{\text{max}})$ continuous value, is the $k=[(x-x_{0})/\Delta]$ where $\cdot$] is used for the integer part (or floor) of the expression. Furthermore, let $y(x)$ a single value continuous function of $x$, i.e. $y:[x_{\text{min}},x_{\text{max}}]\to\mathbb{R}$. Linear interpolation of $y(x)$ at any $x\in[x_{\text{min}},x_{\text{max}})$ such that $x_{k}\leq x<x_{k+1}$ corresponds to
@@
that shows more explicitly that the result of the linear interpolation $\tilde{y}(x)$ is nothing but the linear combination of the two $y_{k},\;y_{k+1}$ exact function values taken at $x_{k}$ and $x_{k+1}$ with the 1w1-w and $w\equiv(x-x_{k})/(x_{k+1}-x_{k})$ weights respectively.
@@
This shows clearly that the result of the linear interpolation of the pdf $\tilde{p}(u;x)$ is nothing but a mixture distribution of the exact $p(u;x_{k})$ and $p(u;x_{k+1})$ pdf-s with the 1w1-w and $w$ weights respectively. This is because the weights 1w1-w and $w$ are both $\geq 0$ while sum up to unity. Therefore, providing $u$ samples distributed according to the $\tilde{p}(u;x)$ interpolated pdf can be done by generating samples form $p(u;x_{k})$ with a probability of 1w1-w and from $p(u;x_{k+1})$ with a probability of $w\equiv(x-x_{k})/(x_{k+1}-x_{k})$.
@@
[…]